### v2

# A Predictive Model for the Dissipation Dilution of Soft-Clamped Modes in Polygon Resonators

Abstract
Polygon resonators are promising candidates for nanomechanical applications due to their compact architecture and high force sensitivity. Here, we develop an analytical framework to predict the resonance frequencies and dissipation dilution factors $D_{Q}$ of polygon perimeter modes by extending the Timoshenko-Gere equation to incorporate the tensile stress. The model identifies two dominant dissipation mechanisms: distributed bending in the polygon sides and torsional deformation in the supporting tethers. We reveal that dissipation dilution in these resonators scales as $1/\lambda^{2}$, distinct from the conventional $1/\lambda$ dependence associated with boundary bending loss. Furthermore, we demonstrate that the torsional loss can be suppressed by tailoring the torsion angle of the supporting tethers. The analytical predictions are validated by finite element simulations, providing a predictive framework for designing high-$Q$ polygon resonators for cavity optomechanics and precision sensing.
Introduction. —
Recent advancements have facilitated ultra-high quality factor ($Q$) nanomechanical resonators, attributed to the phenomenon known as ”dissipation dilution” Tsaturyan (2017); González (2000). This phenomenon originates from the interplay between geometric nonlinearity of strain in deformations and static strain Fedorov et al. (2019). Notably, dissipation dilution significantly minimizes energy loss in the mechanical system through geometric engineering and stress enhancement, rather than relying solely on material optimization Tsaturyan (2017); Fedorov et al. (2019). This opens up exciting avenues for device design.
Investigating the structural design and forces inherent in nanomechanical systems has revealed that the mode curvature at boundaries is a major source of the mechanical energy loss Yu et al. (2012); Tsaturyan (2017). Reducing the amplitude of the mechanical modes at the clamps is an effective way to minimize the energy loss in these systems. Several innovative techniques, referred to as ”soft clamping” Tsaturyan (2017), have been developed specifically for the clamping points to achieve the low-energy-loss. Previous studies have concentrated on tapered clamping Bereyhi et al. (2019), hierarchical structuring Bereyhi et al. (2022b); Fedorov et al. (2020), phononic bandgap engineering Cupertino et al. (2024); Beccari et al. (2022); Ghadimi (2018), and perimeter modes in polygon resonators Bereyhi et al. (2022a).
Remarkably, the perimeter modes in polygon resonators have emerged as promising candidates for integrated quantum optomechanical systems Bereyhi et al. (2022a). In the tapered-clamping approach Bereyhi et al. (2019), enhancement of $Q$ is challenging, as it is constrained to the material’s yield stress-to-deposition stress ratio. In the hierarchical structuring approach Bereyhi et al. (2022b); Fedorov et al. (2020), the torsional losses in cascaded branchings hinder the resonators from attaining the ”clampless” limit. Furthermore, in comparison to the phononic crystals (PnCs) Beccari et al. (2022); Cupertino et al. (2024); Ghadimi (2018), the fundamental perimeter modes of polygon structures Bereyhi et al. (2022a) can reach a high-quality factor of $3.6\times 10^{9}$ with their tenfold reduced size.This compact architecture is advantageous for practical integration with optical microcavities, as it facilitates the maintenance of stable submicron gaps, a prerequisite for achieving strong near-field optomechanical coupling Guo et al. (2019). Additionally, soft clamping in PnCs is typically realized in defect-localized bandgap modes. While bandgap engineering can provide excellent spectral isolation, practical finite-size devices may still host multiple localized defect modes and, depending on boundaries and fabrication disorder, occasional spurious resonances within or near the nominal bandgap. Importantly, soft-clamped perimeter modes in polygon resonators demonstrate exceptional frequency scalability, with a quality factor $Q$ greater than $10^{8}$, maintained across more than four octaves ($170~\text{kHz}-2.5~\text{MHz}$). Moreover, the $Q$s of the perimeter modes exhibit a high thermal-noise-limited force sensitivity of $420~\text{zN}\sqrt{\text{Hz}}$ at room temperature, which is comparable to that of atomic force microscopy cantilevers operating at millikelvin temperature ($190~\text{zN}\sqrt{\text{Hz}}$). Furthermore, the co-integration of perimeter-mode resonators with on-chip optical cavities suppresses both thermomechanical and readout noises, enabling the systematic study of intrinsic frequency fluctuations. The resulting nonlinear optomechanical transduction yields signals with significantly reduced frequency noise, providing a robust platform for precise frequency tracking Kharbanda et al. (2025).
Perimeter modes are advantageous from the perspectives of engineering design and sensing applications. Nevertheless, the current model simulation, which can be matched with experiments and applications, remains relatively preliminary Bereyhi et al. (2022a). Consequently, we develop a comprehensive analytical framework that incorporates tether torsion modes into the existing models and takes into account the torsional resonance in the tethers. In the simulations and analysis, the polygon resonators are composed of a commercial material, silicon nitride (Si3N4). The detailed stress analysis is applied to the Timoshenko-Gere (TG) theory. The model displays remarkable predictive accuracy, with frequency estimations within 0.1% of FEM (finite element method) results across the entire operational range. Furthermore, our findings indicate that boundary losses are substantially amplified by the coupling between the polygon side bending modes and the tether torsional modes. Importantly, these losses can be tailored by adjusting the tether length, providing a flexible strategy for optimizing quality factors.
Figure 1: Schematic representation of the polygon resonator (light blue) with index number $N=4$ and clamps (light blue). The substrates are represented in black. $l_{s}$, $l_{t}$, $w_{s}$, $w_{t}$ represent polygon side length, tether length, polygon side width, and tether width, respectively.
This article is structured as follows. First, we identify two dominant dissipation mechanisms in polygon systems under the narrow-beam approximation. We then derive the vibration of polygon sides, and the supporting tethers, respectively. Consequently, the analytical dissipation dilution can be derived through the modified TG equations by the coupling of these two vibrations. Next, we calculate the $Q$’s of a sample Si3N4 nanomechanical resonator through FEM and compare the results with our analytical results. We can accurately predict the resonance frequencies and the dissipation dilution factors of perimeter modes.
Loss analysis. —
Dissipation dilution can be fundamentally understood through the interaction between a harmonic oscillator and a lossless potential González (2000). The quality factor of a mechanical mode subject to dissipation dilution is given by
$Q=D_{Q}\cdot Q_{\text{int}}.$ (1)
Here, $D_{Q}$ is the dissipation dilution factor, and $Q_{\text{int}}$ is the material’s intrinsic quality factor, which is inversely proportional to the material loss angle ($Q_{\text{int}}=1/\phi$) Saulson (1990). For the $n^{\text{th}}$ resonator mode, the dissipation dilution factor $D_{Q}$ can be written in the general form as Unterreithmeier et al. (2009); Fedorov et al. (2019, 2020):
$D_{\text{Q},n}=\frac{\langle W_{\text{total}}\rangle}{\langle W_{\text{lossy}}\rangle}=\frac{1}{\alpha_{n}\lambda+\beta_{n}\lambda^{2}},$ (2)
where $\langle W_{\text{total}}\rangle$ denotes the total stored energy, and $\langle W_{\text{lossy}}\rangle$ is the lossy part. The coefficients $\alpha_{n}$ and $\beta_{n}$ are determined by the resonator’s boundary curvature and the distributed bending over the rest of the mode, respectively. The strain parameter $\lambda=\frac{h}{l}\sqrt{\frac{E}{12\sigma}}$ depends on the resonator length $l$, thickness $h$, Young’s modulus $E$, and tensile stress $\sigma$.
To study soft-clamped modes in polygon resonators, we start with a four-interconnected beam structure as a general example, as illustrated in Fig. 1. Each beam is supported by two tethers, located at the vertices of the polygon. Due to the underlying symmetry, the fundamental perimeter mode possesses a symmetric displacement profile, as shown in Fig. 2. The transverse reaction forces on opposing tethers demonstrate phase cancellation, resulting in negligible net bending moments. However, the out-of-plane tangential displacement components generate a non-zero torque about the tether axes, inducing torsional dissipation.
Figure 2: FEM simulation of the fundamental mode of a stress preserving polygon resonator. The inset illustrates the torsion deformation caused by the out-of-plane displacement of two neighboring polygon sides. $\tau$ is the torsion angle of the supporting tethers.
In particular, as a result of practical design considerations, the narrow-beam approximation is employed with $l/w\gg 1$ for all constituent beams with length $l$ and width $w$ Bereyhi et al. (2022a). Moreover, previous work has shown that the ratio of the torsional energy stored in a beam to its tensile energy scales as $\lambda^{2}$ Fedorov et al. (2020). This justifies setting the boundary-loss coefficient to be zero ($\alpha_{n}=0$) in the dissipation kernel. Therefore, the dominant dissipation pathways in polygon resonators arise from: (i) bending of the polygonal sides and (ii) shear strain-induced dissipation in the supporting tethers. To analytically estimate the dissipation dilution factor of the perimeter modes, we first derive the equations for the transverse displacement $u(x)$ of the polygon sides and the torsion angle $\tau$ of the supporting tethers, respectively. We then incorporate the coupling between these modes into the model.
Flexural dynamics of polygon sides. —
To model the bending-dominated dissipation, we analyze the transverse displacement field $u(x)$ along each side of the $N$-sided polygon. Within the Euler-Bernoulli beam framework, $u(x)$ satisfies the flexural string equation in the high tension limit Schmid et al. (2016):
$\sigma_{s}(x)\frac{d^{2}u}{dx^{2}}=\rho\omega_{n}^{2}u.$ (3)
where $x$ is the axial coordinate along the polygon sides, $\sigma_{s}$ the stress, $\rho$ the material density, and $\omega_{n}$ the degenerate symmetric frequency of the $n^{\text{th}}$ flexural mode. The stress distribution $\sigma_{s}$ is determined by force balance and the constancy of the path-length between the two clamped vertices of the $N$-sided polygon Bereyhi et al. (2022a):
$\sigma_{s}=\frac{\sigma_{0}(1-\nu)[1+2r_{l}\sin{(\pi/N)}]}{1+4(r_{l}/r_{w})\sin^{2}{}(\pi/N)},$ (4)
with $\sigma_{0}$ representing the initial deposition stress, $\nu$ the Poisson’s ratio, $r_{l}=l_{s}/l_{t}$ the ratio of the polygon side length to the tether length, and $r_{w}=w_{s}/w_{t}$ the ratio of the polygon side width to the tether width. The assumption of perfectly sinusoidal deformation holds under the narrow beam condition ($l\gg w$), reflecting the kinematic constraints of ”soft-clamped” boundaries at vertices. The transverse displacement $u(x)$ can be well approximated by a sinusoidal profile $u(x)=u_{0}\sin{(n\pi x/l_{s})}$ with polygon side length $l_{s}$. The degenerate symmetric frequency $\omega_{n}$ can be determined by substituting $u(x)$ into Eq. 3:
$\omega_{n}^{2}=\frac{\sigma_{s}}{\rho}\frac{n^{2}\pi^{2}}{l_{s}^{2}}.$ (5)
Torsional vibration of supporting tethers. —
The torsion angle $\tau$ is determined by the modified TG equation, incorporating the axial stress term (see Appendix A for the full derivation of the modified TG equation):
$EI_{\psi}\frac{\partial^{4}\tau}{\partial x^{4}}+(\rho I_{\psi}\omega_{n}^{2}-C-\sigma_{t}I_{y})\frac{\partial^{2}\tau}{\partial x^{2}}-\rho I_{p}\omega_{n}^{2}\tau=0.$ (6)
where $\sigma_{t}=\sigma_{xx}$ denotes the axial stress along the tether, $x,y,z$ are the cartesian coordinates, $I_{p}=\iint_{A}(y^{2}+z^{2})\mathrm{d}A$ is the geometric moment of inertia, $I_{\psi}=\iint_{A}\psi^{2}\mathrm{d}A$ ($\psi$ $\approx-yz$ the warping function Chopin and Filho (2019)), $A$ the cross section with thickness $h$ and tether width $w_{t}$, $C$ the torsional rigidity ($C=\iint_{A}G[(\frac{\partial\psi}{\partial y}-z)^{2}+(\frac{\partial\psi}{\partial z}+y)^{2}]\mathrm{d}A$), and $G$ the shear modulus. We note that Eq. 6 is structurally analogous to the Euler-Bernoulli beam equation in both the structure and mode shape representation (Schmid et al. (2016) (p46), and Eq. S5). Consequently, the approximation method, used for high-stress conditions in Eq. 3, is equally applicable to Eq. 6, which can be simplified to
$(C+\sigma_{t}I_{y})\frac{d^{2}\tau}{dx^{2}}=\rho I_{\text{p}}\omega_{n}^{2}\tau,$ (7)
where the axial stress $\sigma_{t}$ represents the enhanced effective torsional rigidity of the tether through the term $C+\sigma_{t}I_{y}$. To solve Eq. 7, we assume a trial solution $\tau(x)=\tau_{0}\sin{(kx+\phi_{0})}$ with the boundary condition $\tau(0)=0$, therefore the wave number $k$ is identified as
$k=\sqrt{I_{\text{p}}\sigma_{s}\frac{n^{2}\pi^{2}}{l_{s}^{2}}\frac{1}{C+\sigma_{t}I_{y}}}.$ (8)
Coupling at the branch nodes. —
The flexural deformation of the polygon sides generates a torque that is directly coupled to the supporting tethers, initiating torsional wave propagation along the tether elements. The coupling between the transverse and torsional vibrations can be characterized by the following relation Fedorov et al. (2020):
$u^{\prime}(0)=\sin[\frac{(N-2)\pi}{2N}]\tau(l_{t})=\cos{(\frac{\pi}{N})}\tau_{0}\sin{(kl_{t})}.$ (9)
Eq. 9 demonstrates that the modal gradient at branch nodes, $u^{\prime}(0)$, is determined by both the cosine of the polygon side number $N$ and $\sin{(kl_{t})}$. Since the system’s tensile energy scales quadratically with $u^{\prime}(0)$, theoretical $Q$ maximization requires: (i) increasing the number of polygon sides, and (ii) achieving $\sin{(kl_{t})}=1$. However, the former has been shown to offer a limited practical improvement in dissipation dilution Bereyhi et al. (2022b); Fedorov et al. (2020). Importantly, the latter provides a novel insight: when all material parameters except $l_{t}$ are fixed (wave number $k=\text{const}$), the modal gradient $u^{\prime}(0)$ - and consequently the dissipation dilution factor $D_{Q}$ - can be optimized by adjusting $l_{t}$.
Dissipation dilution factors and loss coefficient. —
The dissipation dilution factor $D_{Q}$ of perimeter modes can be determined by estimating different energy components in Eq. 2. The ”lossless” tensile energy is given by Fedorov et al. (2020)
$\langle W_{\text{tens}}\rangle=\frac{N}{2}\sigma_{s}w_{s}h\int_{0}^{l_{s}}[u^{\prime}(x)]^{2}dx.$ (10)
The lossy energy comprises two components. The distributed bending energy along the polygon sides is
$\langle W_{\text{bend}}\rangle=\frac{N}{2}\frac{Ew_{s}h^{3}}{12}\int_{0}^{l_{s}}[u^{{}^{\prime\prime}}(x)]^{2}dx,$ (11)
and the torsional lossy part is
$\langle W_{\text{tors}}\rangle=N\frac{Gw_{t}h^{3}}{12}\int_{0}^{l_{t}}(\tau^{\prime}(x))^{2}dx.$ (12)
Substituting Eqs. 10 - 12 into Eq. 2 and defining $\lambda_{l}=\frac{h}{l_{s}}\sqrt{\frac{E}{12\sigma_{s}}}$ yields a dissipation dilution factor of
$\begin{split}D_{Q,n}^{-1}&=\frac{\langle W_{\text{bend}}\rangle+\langle W_{\text{tors}}\rangle}{\langle W_{\text{tens}}\rangle}\\ &=D_{Q,\text{bend}}^{-1}+D_{Q,\text{tors}}^{-1}.\end{split}$ (13)
with the two lossy contributions
$D^{-1}_{Q,\text{bend}}=n^{2}\pi^{2}\lambda_{l}^{2},$ (14)
and
$D^{-1}_{Q,\text{tors}}=\lambda_{l}^{2}\frac{kl_{s}}{(1+\nu)r_{w}}\frac{2kl_{t}+\sin{2kl_{t}}}{\cos^{2}(\pi/N)\sin^{2}(kl_{t})}.$ (15)
Figure 3: Dissipation dilution factors and frequencies of the perimeter modes of the resonator with dimensions shown in the simulation part. Blue filled circles and red open circles correspond to FEM results and analytical results, respectively. Green squares correspond to out-of-plane modes of normal strings with the same dimensions and mechanical properties as the polygon sides. The four insets show the fundamental perimeter mode (a), and higher order perimeter modes where bending-torsion coupling exists (b, c, and d).
The first term in Eq. 13 is due to the distributed bending across the polygon sides, and sets an upper bound on the dissipation dilution factor. The second term is related to the torsional deformation of the supporting tethers. Notably, the term $\sin^{2}{(kl_{t})}$ in Eq. 15 implies that the dissipation dilution factor $D_{Q,n}$ has a lower bound 0 when $kl_{t}$ is an integer multiple of $\pi$. In this case, the polygon sides exhibit zero transverse displacement, resulting in a system where the total energy manifests exclusively as dissipative torsional energy. To achieve the optimal dilution factor, $l_{t}$ should be chosen to ensure $kl_{t}=\pi/2$. It is crucial that the torsional loss part $D_{Q,\text{tors}}$ depends inversely on $\lambda_{l}^{2}$, thereby establishing a scaling law fundamentally different from the $1/\lambda$ dependence of clamping loss in conventional dissipation dilution Fedorov et al. (2020); Bereyhi et al. (2022b, 2019). This insight promises a new paradigm for engineering ultra-high-Q resonators, one that circumvents the need for the large dimensions historically required by phononic crystals to open a bandgap and approach the ”clampless” limit Ghadimi (2018).
Simulation. —
As an illustrative case, we consider a polygon resonator characterized by the following geometric parameters: vertex count $N=4$, side length $l_{s}=1400~\upmu\mathrm{m}$, polygon side width $w_{s}=300~\mathrm{nm}$, support length $l_{t}=700~\upmu\mathrm{m}$, and support width $w_{t}=400~\mathrm{nm}$. The resonator structure is made of a $100~\mathrm{nm}$ commercial Si3N4 thin film with Young’s modulus $E=250~\mathrm{GPa}$, mass density $\rho=3100~\mathrm{kg/m^{3}}$, Poisson’s ratio $\nu=0.23$, and deposition stress $\sigma_{0}=1~\mathrm{GPa}$.
Figure 4: Dissipation dilution factors of fundamental perimeter modes with $N=4$. (a) $D_{Q}$ dependence on the tether length, with $l_{s}=700~\upmu\mathrm{m}$, $w_{s}=300~\mathrm{nm}$, and $w_{t}=400~\mathrm{nm}$. The orange triangles and light blue circles represent the FEM simulation results and the analytical predictions, respectively. (b) Torsion angle profiles for the tether at lengths of 3920, 7890, and $11810~\upmu\mathrm{m}$. The results from the FEM simulation are shown in orange and are compared against sinusoidal references (light blue). (c) $D_{Q}$ dependence on the polygon width, with $l_{s}=700~\upmu\mathrm{m}$, $l_{t}=700~\upmu\mathrm{m}$, and $w_{t}=400~\mathrm{nm}$. (d) Distribution of $D_{Q}$ for $l_{s}=700~\upmu\mathrm{m}$ and $w_{s}=300~\mathrm{nm}$. (e) Distribution of $D_{Q}$ for $l_{t}=700~\upmu\mathrm{m}$ and $w_{t}=400~\mathrm{nm}$.
[…]
To demonstrate the precision of the analytical model, we present in Fig. 3 a comparative analysis of resonance frequencies and dissipation dilution factors for 20 modes of the polygon resonator, as obtained through both the FEM and analytical modeling. The analytical model predicts the resonance frequencies with relative errors less than 0.1%. For most modes, the analytically estimated dilution factors are in good quantitative agreement with the FEM results. However, anomalies are observed in the $7^{\mathrm{th}}$ and $14^{\mathrm{th}}$ perimeter modes due to internal resonance effects, where energy transfers from primary flexural modes to coupled torsional modes (comparison between Insets a, b, and d reveals distinct torsional features). This energy transfer leads to significant $Q$-factor degradation in primary modes Asadi et al. (2021). In addition, the $11^{\mathrm{th}}$ and $12^{\mathrm{th}}$ modes (inset c) exhibit deviation from theoretical predictions due to the coupling of the bending-torsion mode in the tethers. Remarkably, the dilution factor of the fundamental mode is approximately two orders of magnitude higher than that of a simple doubly-clamped beam whose dimensions match those of a single polygon side.
To investigate the hybridization between the bending modes of the polygon sides and the torsional modes of the tethers, we fix $l_{\text{s}}=700~\upmu\mathrm{m}$ and vary the length of the tethers while keeping all other parameters constant. Fig. 4 plots the dissipation dilution factor as a function of the tether length in the polygon resonator, revealing three well-defined resonant peaks. The first peak yields the maximum dissipation dilution, while subsequent peaks exhibit progressively reduced magnitudes. This behavior originates from the torsional rotation at the tether-polygon junction, which modulates the slope of the transverse displacement of the polygon side - a quantity whose squared value directly governs the tensile energy based on Eq. 9. Consistent with the model, both the first peak at $kl_{t}\approx\pi/2$ and the first trough at $kl_{t}\approx\pi$ are in excellent agreement with Eq. 15. The torsion-angle profiles at $kl_{t}\approx\pi$, which correspond to tether lengths of 3920, 7890, and $11810~\upmu\mathrm{m}$, are further visualized in Fig. 4. These results indicate that adjustment of the tether length allows precise control over the quality factors of the resonator and offers valuable guidance to optimize high-performance polygon resonators. As the tether length increases, we also notice that the peak positions predicted by the analytical model exhibit a growing deviation from the FEM simulations. This physical mechanism underlying the increase in $k$ with $l_{\text{t}}$ remains under investigation.
While our theoretical model accounts for distributed bending loss in the polygon sides and torsional loss in the tethers, it neglects the torsional loss in the polygon sides; this simplification could also cause the discrepancy between the analytical model and FEM simulation, especially for wider designs. Analogous to neglecting the boundary loss of sufficiently long polygon sides, we can also define a range for the polygon side width $w_{s}$ within which its torsional loss is negligible. To validate this, we perform simulations with $l_{s}=l_{t}=700~\upmu\mathrm{m}$ and $w_{t}=400~\mathrm{nm}$, and scan $w_{s}$ from $100~\mathrm{nm}$ to $2000~\mathrm{nm}$. As shown in Fig. 4, the dissipation dilution factor increases monotonically as the width increases from $100~\mathrm{nm}$ to $400~\mathrm{nm}$, suggesting that the torsional loss is negligible in this regime. For widths from $400~\mathrm{nm}$ to $2000~\mathrm{nm}$, the close agreement between our theoretical predictions and the FEM results, with a deviation of merely 6%, further confirms that the polygon side torsional loss has an insignificant impact across the entire studied range.
To systematically explore the impact of resonator dimensions on the quality factor and identify the optimal configuration, we employ a parametric approach, by independently varying tether and polygon side dimensions. First, we fix $l_{s}=700~\upmu\mathrm{m}$, $w_{s}=300~\mathrm{nm}$, while scanning the tether length $l_{t}$ over a range of $1-4000~\upmu\mathrm{m}$ and the tether width $w_{t}$ over $1-2000~\mathrm{nm}$. The results, as shown in Fig. 4, are consistent with those in Fig. 4. Several local maxima of $D_{Q}$ are observed, with the highest value occurring at point ($1700~\upmu\mathrm{m}$, $450~\mathrm{nm}$). In Fig. 4, with the tether length and width fixed at $700~\upmu\mathrm{m}$ and $400~\mathrm{nm}$, respectively, $D_{Q}$ increases monotonically with $l_{s}$. This is primarily driven by the $D_{Q}\propto l_{s}^{2}$ dependence originating from the $\lambda_{l}^{2}$ term in Eq. 14 and Eq. 15. Moreover, the optimal $w_{s}$ increases with $l_{s}$, which can be attributed to the linear dependence of the $kl_{s}/[(1+\nu)r_{w}]$ term in Eq. 15 on $l_{s}$.
Our analytical model provides a reliable estimate of the dissipation dilution factor $D_{Q}$ for perimeter modes, accurately capturing its dependence on the mode number, polygon side width, and tether length. As shown in Fig. 4, the model agrees well with FEM simulations for short and moderate long tethers. The noticeable discrepancy emerges in the limit of extremely long tethers or excessively wide side widths, which can be attributed to a simplified assumption in the model, namely that the axial stress in the tethers remains constant while the tether length varies. Prior studies Bückle et al. (2021) have indicated that increasing the tether length reduces the axial stress. Within our model, this stress relaxation modifies the effective torsional stiffness, as described in Eq. 7, leading to a weak dependence of $k$ on $l_{t}$ (Eq. 8) and a corresponding shift in the resonance peaks. Meanwhile, numerical convergence challenges in FEM for extremely high-aspect-ratio geometries cannot be excluded, although they are not expected to dominate the noticeable discrepancy. It is worth noting that, while the effect of the torsional loss in the polygon sides is neglected here, the side width corresponding to the maximum $D_{Q}$ for a fixed polygon side length lies within a range where the torsional effect in the polygon sides is negligible.
Torsional dissipation dilution limit. —
An important issue in the study of polygon structures involves the enhancement of dissipation dilution in the supporting tethers. This phenomenon has been demonstrated in doubly-clamped stressed nanoribbons, which are analogous to the supporting tethers in polygon structures (see Appendix B for a detailed derivation). Thus, considering the tethers within a polygon resonator system, the dissipation dilution factor for the $n^{\text{th}}$ torsional mode is given by:
$D_{Q,t,n}\approx 1+\frac{1}{\frac{2G}{\sigma_{t}}\frac{h^{2}}{w_{t}^{2}}+\frac{n^{2}\pi^{2}}{12}\frac{Eh^{2}}{\sigma_{t}l_{t}^{2}}+\frac{h}{l_{t}}\sqrt{\frac{E}{3\sigma_{t}}}}.$ (16)
The dimensions of the tethers ($l_{t}\gg w_{t}$) and material properties of the Si3N4 film (see the simulation part) suggest that the first term $\frac{2G}{\sigma_{t}}\frac{h^{2}}{w_{t}^{2}}$ in the denominator of Eq. 16 dominates the energy dissipation. This finding aligns with prior work Pratt et al. (2023) and points to a limitation in torsional dissipation dilution.
Conclusions and outlook. —
In this work, we present a comprehensive analytical study of perimeter vibrational modes in polygonal resonators. We derive the TG equation for the angular displacement of nanoribbons under the high tension condition. The modified framework can be extended to the torsional vibration of supporting tethers of polygon resonators, accurately predicting both the resonance frequencies and dissipation dilution factors of perimeter modes. Notably, the quality factor of the fundamental mode is enhanced by two orders of magnitude compared to uniform strings of the same dimensions as the polygon sides. Our analysis reveals that the resonance of the tethers can significantly degrade the dissipation dilution factor of the perimeter mode, and the tether length can be used to tune the dilution factor. Our findings reveal that the dissipation dilution factor of the perimeter mode in polygon resonators scales as $1/\lambda^{2}$, consistent with the existing literature Bereyhi et al. (2022a). This scaling law provides a transformative strategy for designing ultra-high-$Q$ resonators, with direct applications in cavity optomechanics Aspelmeyer et al. (2014). Concurrently, we elucidate the torsional dissipation dilution limit of the supporting tethers. Additionally, our work has significant implications for nanoscale force sensing Moser et al. (2013); Gavartin et al. (2012), enabling precision measurements in weak-force detection. Furthermore, our work advances the study of multimode phononic structures Faust et al. (2012); Peano et al. (2015); Lu et al. (2017), offering new avenues for exploring topological phononics and non-Hermitian wave control.
I Acknowledgements
Y. Z. acknowledges the support from the National Natural Science Foundation of China (Grant No. 12474153, Grant No. 11804163), Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401001).
Appendix A Appendix A: Free Torsional Vibration Under Tensile Stress
The theoretical foundation of torsional dissipation dilution can be traced back to Buckley’s pioneering work on twisted elastic strips Buckley (1914), which demonstrates that tensile stress can geometrically enhance torsional stiffness without proportionally increasing dissipative losses. This phenomenon arises because tensile stress preferentially aligns the principal strain axes along the ribbon’s length, thereby suppressing transverse deformation through geometric stiffening, while intrinsic loss mechanisms remain primarily governed by material properties Trahair (2017) (p258-265). Buckley’s analysis of bifilar systems reveals that torsion modes of ribbons are naturally soft-clamped Tsaturyan (2017) and the effective quality factor $Q_{\mathrm{eff}}$ can be enhanced beyond the material limit through stress engineering, a critical insight for modern nanomechanical resonators Fedorov et al. (2019).
The study by Pratt et al. Pratt et al. (2023) significantly extends Buckley’s paradigm by demonstrating experimentally the validation of extreme dissipation dilution ($Q\cdot f>10^{13}$ Hz) in nanoribbons made of $\text{Si}_{3}\text{N}_{4}$ at room temperatures. Their work establishes a novel theoretical framework for quantum precision measurements exploiting the torsion modes of nanomechanical resonators Kim et al. (2016); Enomoto et al. (2016); Treps et al. (2003), enabling novel designs where torsional modes benefit from dissipation dilution Bereyhi et al. (2022a); Ghadimi (2018).
Here, we derive the general solution for the torsional vibrational modes of a stressed nanoribbon based on Hamilton’s variational principle. First, we revisit the TG theory for torsional vibration of ribbons in the cartesian representation Rao (2019) (p296-302). For a normal nanoribbon resonator with length $L$, width $w$, and thickness $h$ satisfying $L>w>h$, the equation of motion for the angular displacement $\Theta(x,t)$ is
$\begin{split}\rho I_{p}\frac{\partial^{2}\Theta}{\partial t^{2}}-\frac{\partial^{2}}{\partial x\partial t}(\rho I_{\psi}\frac{\partial^{2}\Theta}{\partial x\partial t})&-\frac{\partial}{\partial x}(C\frac{\partial\Theta}{\partial x})\\ &+\frac{\partial^{2}}{\partial x^{2}}(EI_{\psi}\frac{\partial^{2}\Theta}{\partial x^{2}})=0,\end{split}$ (A1)
where $x$ denotes the axial coordinate along the ribbon, $I_{p}=\iint_{A}(y^{2}+z^{2})\mathrm{d}A$ is the geometric moment of inertia, $I_{\psi}=\iint_{A}\psi^{2}\mathrm{d}A$ ($\psi$ $\approx-yz$ is the warping function Chopin and Filho (2019)), $A=w\cdot h$ the cross section, $C$ the torsional rigidity ($C=\iint_{A}G[(\frac{\partial\psi}{\partial y}-z)^{2}+(\frac{\partial\psi}{\partial z}+y)^{2}]\mathrm{d}A$), $E$ the Young’s modulus, $G$ the shear modulus, and $\rho$ the material density. In the case of a highly stressed ribbon, the angular displacement is obtained by minimizing the action integral $\mathcal{S}=\int_{t_{1}}^{t_{2}}(T-W_{\text{strain}}-W_{\text{tens}})\,dt$, where $T$, $W_{\text{strain}}$ and $W_{\text{tens}}$ represent the kinetic, elastic strain and tensile energies, respectively. The modified TG equation can be written as
$\begin{split}\rho I_{p}\frac{\partial^{2}\Theta}{\partial t^{2}}-\frac{\partial^{2}}{\partial x\partial t}(\rho I_{\psi}\frac{\partial^{2}\Theta}{\partial x\partial t})&-\frac{\partial}{\partial x}((C+\sigma I_{y})\frac{\partial\Theta}{\partial x})\\ &+\frac{\partial^{2}}{\partial x^{2}}(EI_{\psi}\frac{\partial^{2}\Theta}{\partial x^{2}})=0,\end{split}$ (A2)
where $I_{y}$ is the geometric moment of inertia with respect to the $y$-axis, and $\sigma=\sigma_{xx}$ is the axial stress component. By applying separation of variables, the solution to this differential equation takes the form of a superposition of normal modes, each consisting of a position-dependent component multiplied by a time-dependent component:
$\Theta(x,t)=\sum_{n=1}^{\infty}u_{0,n}\cos{(\omega_{n}t)}\theta_{n},$ (A3)
where $\omega_{n}$ is the eigenfrequency of the $n^{\text{th}}$ mode, $\theta_{n}$ the mode shape, and $u_{0,n}$ the amplitude. Inserting Eq. A3 into Eq. A2 yields
$EI_{\psi}\frac{\partial^{4}\theta_{n}}{\partial x^{4}}+(\rho I_{\psi}\omega_{n}^{2}-C-\sigma I_{y})\frac{\partial^{2}\theta_{n}}{\partial x^{2}}-\rho I_{p}\omega_{n}^{2}\theta_{n}=0.$ (A4)
Assuming a sinusoidal mode shape with wavenumber $\beta_{n}$, and applying the boundary conditions $\theta_{n}(0)=\theta_{n}(L)=\frac{\partial^{2}\theta_{n}}{\partial x^{2}}|_{0}=\frac{\partial^{2}\theta_{n}}{\partial x^{2}}|_{L}=0$ yield
$\theta_{n}(x)=\sin{(\beta_{n}x)},~\beta_{n}=\frac{n\pi}{L}.$ (A5)
Combining Eq. A5 with Eq. A4 gives
$\begin{split}\rho w^{2}h^{2}\beta_{n}^{2}\omega_{n}^{2}+12\rho(w^{2}+h^{2})\omega_{n}^{2}&=Ew^{2}h^{2}\beta_{n}^{4}\\ &+48Gh^{2}\beta_{n}^{2}+12w^{2}\sigma\beta_{n}^{2}.\end{split}$ (A6)
The first term $\rho w^{2}h^{2}\beta_{n}^{2}\omega_{n}^{2}$ is negligible due to $\rho w^{2}h^{2}\beta_{n}^{2}\ll 12\rho(w^{2}+h^{2})$. Similarly, the term $Ew^{2}h^{2}\beta_{n}^{4}$ exhibits negligible influence on the system dynamics compared to the two dominant terms dominating the right-hand side. The mode frequency $\omega_{n}$ can be written in the following form
$\begin{split}\omega_{n}&=\sqrt{\frac{48Gh^{2}\beta_{n}^{2}+12w^{2}\sigma\beta_{n}^{2}}{12\rho(w^{2}+h^{2})}}\\ &\approx\frac{n\pi}{L}\sqrt{\frac{\sigma}{\rho}(1+\frac{4G}{\sigma}\frac{h^{2}}{w^{2}})}.\end{split}$ (A7)
The full set of wave numbers can be found by solving the dispersion relation of the pre-stressed ribbon:
$\beta_{n}^{4}+(\frac{48G}{Ew^{2}}+\frac{12\sigma}{Eh^{2}})\beta_{n}^{2}-\frac{12\rho}{Eh^{2}}\omega_{n}^{2}=0.$ (A8)
Four solutions of $\beta_{n}$ can be found:
$\beta_{n,1-2}=\beta_{\sigma}=\frac{n\pi}{L},$ (A9)
$\beta_{n,3-4}=\beta_{E}=\sqrt{\frac{12\sigma}{Eh^{2}}(1+\frac{4G}{\sigma}\frac{h^{2}}{w^{2}})}.$ (A10)
Within the analytical model, $\beta_{\sigma}$ is the already known wavenumber of a perfect ribbon, and $\beta_{E}$ is the wave number related to the torsional stiffness of the ribbon. Under clamped boundary conditions, the mode shape takes the following form:
$\theta_{n}(x)=\sin{(\beta_{\sigma}x)}+\frac{\beta_{\sigma}}{\beta_{E}}[e^{-\beta_{E}x}-\cos{(\beta_{\sigma}x)}].$ (A11)
This equation enables the decoupling of the idealized torsion mode from the boundary-constrained components in the actual vibrational mode. It reveals that $\beta_{E}$ corresponds to the inverse characteristic length of the exponentially decaying edge effect.
Appendix B Appendix B: Torsional Dissipation Dilution of Nanoribbons
The quality factor of a stressed nanomechanical resonator exhibits significant enhancement due to dissipation dilution. For a ribbon resonator under tensile stress $\sigma$, the dissipation dilution factor $D_{Q}$ can be written as
$D_{Q,n}=\frac{Q}{Q_{\text{int}}}=1+\frac{\langle W_{\text{tens}}\rangle}{\langle W_{\text{lossy}}\rangle}$}, (B1)
where $\langle W_{\text{tens}}\rangle$ is the energy stored in the stress potential, and $\langle W_{\text{lossy}}\rangle$ is the energy concentrated in the bending and shear components Fedorov et al. (2019); Sementilli et al. (2022). The dissipation dilution factor $D_{Q}$ measures the proportional enhancement in the mechanical quality factor resulting from tensile stress. To analytically estimate $D_{Q}$, we express a lossless tensile energy based on Eq. A3:
$\langle W_{\text{tens}}\rangle=\frac{\sigma}{2}\frac{hw^{3}}{12}\int u_{0,n}^{2}\langle(\frac{\partial\theta_{n}}{\partial x})^{2}\rangle dx.$ (B2)
The lossy energy is
$\langle W_{\text{lossy}}\rangle=\frac{E}{2}\frac{wh^{3}}{12}\int u_{0,n}^{2}\langle\frac{2}{1+\nu}(\frac{\partial\theta_{n}}{\partial x})^{2}+\frac{w^{2}}{12}(\frac{\partial^{2}\theta_{n}}{\partial x^{2}})^{2}\rangle dx.$ (B3)
Substituting Eq. B2 and Eq. B3 into Eq. B1 yields a dissipation dilution factor of
$D_{\text{Q},n}=1+\frac{1}{\frac{2G}{\sigma}\frac{h^{2}}{w^{2}}+\frac{n^{2}\pi^{2}}{12}\frac{Eh^{2}}{\sigma L^{2}}+\frac{h}{L}\sqrt{\frac{E}{3\sigma}(1+\frac{4G}{\sigma}\frac{h^{2}}{w^{2}})}}.$ (B4)
The first two terms in the denominator arise from distributed shear deformation and distributed bending, respectively, while the last term is due to the coupled bending-shear interaction at the clamps. For a standard long-string resonator, the length is much greater than the thickness, causing the first term to dominate Sadeghi et al. (2020). In commercial high-stress silicon nitride, the ratio of Young’s modulus to stress is approximately 250 Fedorov et al. (2020), leading to a calculated $D_{Q}$ of $\sim 1$. Consequently, the torsion mode cannot experience dissipation dilution Fedorov et al. (2019). In contrast, for a nanoribbon, the width substantially exceeds the thickness as well, resulting in Q factors which scale as $(w/h)^{2}$ as high as $10^{8}$ Pratt et al. (2023). Provided that curvature at the clamping points is nonnegligible, Eq. B4 can be reduced to
$D_{Q,n}\approx 1+\frac{1}{\frac{2G}{\sigma}\frac{h^{2}}{w^{2}}+\frac{n^{2}\pi^{2}}{12}\frac{Eh^{2}}{\sigma L^{2}}+\frac{h}{L}\sqrt{\frac{E}{3\sigma}}}.$ (B5)
The formulation bears strong similarity to the dissipation dilution factor expression for standard string resonators Schmid et al. (2016); Sadeghi et al. (2019). This correspondence suggests that established techniques for enhancing the quality factor in flexural modes of string resonators - including but not limited to phononic crystal engineering Beccari et al. (2022), tapered clamping structures Bereyhi et al. (2019), and hierarchical mechanical designs Bereyhi et al. (2022b); Fedorov et al. (2020)- can be effectively extended to enhance the torsion mode performance in ribbon resonators.